How many cubic centimeters can a cigar box hold if its dimensions are 3 centimeters by 16 centimeters by 12 centimeters?
Answer by jim_thompson5910(35256)

Volume = L*W*H = 3*16*12 = 3*192 = 576


So the volume of the box is 576 cubic centimeters